OpenStudy (anonymous):

how do i solve 256=r^4

OpenStudy (anonymous):

Find the fourth root of both sides: \[\sqrt[4]{256} = \sqrt[4]{r}^{4}\] The fourth root of 256 is 4, therefore r = 4

OpenStudy (anonymous):

is there a calculator function for that?

OpenStudy (anonymous):

or should i just guess to find answer?

OpenStudy (anonymous):

There should be, but you could also put it to the power of 1/4, for example: \[256^{1/4}\]

OpenStudy (anonymous):

Some calculators have it, and others do not - but putting the answer to the power to 1/4 should always work. If you were finding a fifth root, you could do the same thing, except you would use the power of 1/5.

OpenStudy (anonymous):

i see when i type it in i get 64 though

OpenStudy (anonymous):

is my calculator being weird? your math makes sense

OpenStudy (anonymous):

never mind i figured it out. thank you though

OpenStudy (anonymous):

You could try this calculator: http://web2.0calc.com/ I just used it and it gave me 4. Make sure you're using the exponent button and not accidentally hitting anything else

OpenStudy (anonymous):

No problem!

OpenStudy (anonymous):

for anyone who has a ti-83 plus and has a simmilar problem Example: Find the 5th root of 16807. • Input 5. • Press [MATH] [5] to input the nth root command. • Input 16807 and press [ENTER] to complete the calculation.

OpenStudy (raden):

looks the possibly other is -4 (-4)^4 = 256 (correct too)

OpenStudy (raden):

r^4 = 256 r^4 - 256 = 0 (r^2)^2 - 16^2 = 0 (r^2 + 16)(r^2 - 16) = 0 r^2 + 16 = 0 (no roots are real) r^2 - 16 = 0 (r+4)(r-4) = 0 the real roots satisfied for r = 4 or r = -4